# Generating Model Parameters for Controlling: Parameter Diffusion for Controllable Multi-Task Recommendation

### v1

@@ Abstract
Commercial recommender systems face the challenge that task requirements from platforms or users often change dynamically (e.g., varying preferences for accuracy or diversity). Ideally, the model should be re-trained after resetting a new objective function, adapting to these changes in task requirements. However, in practice, the high computational costs associated with retraining make this process impractical for models already deployed to online environments. This raises a new challenging problem: how to efficiently adapt the learning model to different task requirements by controlling model parameters after deployment, without the need for retraining. To address this issue, we propose a novel controllable learning approach via Parameter Diffusion for controllable multi-task Recommendation (PaDiRec), which allows the customization and adaptation of recommendation model parameters to new task requirements without retraining. Specifically, we first obtain the optimized model parameters through adapter tunning based on the feasible task requirements. Then, we utilize the diffusion model as a parameter generator, employing classifier-free guidance in conditional training to learn the distribution of optimized model parameters under various task requirements. Finally, the diffusion model is applied to effectively generate model parameters in a test-time adaptation manner given task requirements. As a model-agnostic approach, PaDiRec can leverage existing recommendation models as backbones to enhance their controllability. Extensive experiments on public datasets and a dataset from a commercial app, indicate that PaDiRec can effectively enhance controllability through efficient model parameter generation. The code is released at https://anonymous.4open.science/r/PaDiRec-DD13e.
22footnotetext: Corresponding author.11footnotetext: These authors contributed equally.
1 Introduction
Traditional recommender systems are usually designed to improve accuracy by analyzing user behaviors and contextual data to uncover users’ potential interests and preferences (Kang & McAuley, 2018; Hidasi et al., 2016). Nowadays, recommendation models place greater emphasis on multiple important aspects of the recommended results (also called multi-task recommendation), such as diversity (Xia et al., 2017), fairness (Oosterhuis, 2021), etc. Existing multi-task recommendation models are typically static (Zhang & Yang, 2021; Sener & Koltun, 2018), meaning that the preference weights for each aspect (e.g., accuracy or diversity) are predefined and fixed during both training and testing. Once the static preference weights are determined, the training process can employ various optimization algorithms to find the optimal solution.
However, in practical scenarios, the preference weights for different aspects often change dynamically across both scenarios and time (Hariri et al., 2015; Shen et al., 2023). From a commercial perspective, different application scenarios may require varying preference weights for different performance aspects of the recommendation model to meet specific business needs. For instance, the checkout page emphasizes product diversity, while the product detail page prioritizes accuracy by recommending similar items. From the users’ perspective, different user groups may have distinct preferences, and even the same users may have changing information needs over time. For example, a user may prefer highly accurate recommendations when browsing a specific item category, but over time, such precision might diminish their interest, prompting a preference for more diverse categories. To address the above dynamic information needs of users or platforms, this paper focuses on enhancing the controllability of recommendation models at test time, specifically in the context of controllable multi-task recommendation.
Traditional multi-task learning approaches face challenges in addressing the issue of dynamically changing preference weights. More specifically, when preference weights change, they require resetting the objective function, re-training the recommendation model based on the new objective, and then redeploying the updated model. However, while this approach enables the integration of various optimization methods, the retraining process is highly time- and resource-intensive, rendering it impractical — especially since rapid response time is critical during online recommendation phases. For instance, during promotional events, commercial stores often require real-time flow control to adjust their recommendation strategies, with changes ideally implemented immediately. Several studies have recognized the importance of dynamically adjusting models based on changing preferences. Wortsman et al. (2022) used simple parameter merging across multiple task-specific models, and Chen et al. (2023) employed discriminative models to generate parameters for multi-task re-ranking problem. While they reduce response time and aim to enhance control over the model, they struggle with approximating the optimal model (which we assume can be achieved through retraining with given preference weights), potentially leading to suboptimal solutions. To achieve both efficient test-time adaptation to changing preferences and preserve the approximate optimal performance that retraining offers, we leverage the strengths of diffusion models in generating high-performance model parameters (Schürholt et al., 2022; Knyazev et al., 2021; Wang et al., 2024) for recommendation model. Additionally, we utilize conditional control (Ho & Salimans, 2022) to ensure controllability at test-time with changing preference weights as conditions.
In this work, we propose a novel parameter generation approach for controllable multi-task recommendation by leveraging a generative model to efficiently generate task-specific model parameters at test time based on varying task requirements (i.e., the preference weights for different performance metrics), effectively addressing the challenges posed by rapidly changing requirements and the high cost of retraining models. The proposed approach, termed PaDiRec, begins by formulating an objective function aligned with task-specific preference weights, and through advanced optimization techniques, we fine-tune model parameters using adapter tuning. We then train a diffusion model to learn the conditional distribution of these optimized adapter parameters under various task requirements, where the classfier-free guidance training strategy is employed to perform conditional training. Once trained, during online testing, the diffusion model can generate task-specific adapter parameters with the task requirement as condition, which can be integrated with different sequential recommendation backbones to produce recommendation lists that meet the specified requirements. Additionally, PaDiRec is both model-agnostic and algorithm-agnostic, making it flexible and compatible with various recommendation models and optimization strategies. We summarize our contributions as follows:
  • We formally define the problem of controllable multi-task recommendation (CMTR), which focuses on the model’s ability to adapt to dynamic changes in preferences for different metrics during online testing.
  • We present PaDiRec, a diffusion model-based approach that generates model parameters conditioned on task-specific preference weights, providing enhanced control and flexibility by controlling model parameters in multi-task learning settings.
  • Extensive experiments on two public datasets and an industrial dataset demonstrate that PaDiRec achieves superior performance towards controllability of multi-task recommendation while retaining recommendation performances.
2 Problem Formulation and Analyses
@@
To efficiently and effectively adapt to changes in tasks, this paper focuses on controlling the model parameters $\bm{\theta}$ of the recommendation model $f_{\bm{\theta}}$ to accommodate the varying preference weights of new tasks. More specifically, we treat the preference weights as variables and model the relationship between the preference weight space $\mathcal{W}$ and the model parameter space $\Theta$ during training, transforming the time- and resource-intensive retraining problem at test time into an efficient inference problem. Formally, we aim to find a function $g_{\bm{\xi}}:\mathcal{W}\rightarrow\Theta$ (where $\bm{\xi}$ denotes the parameter of $g$) that generates model parameters capable of achieving a high reward given the new preference weights $\bm{w}_{k}$ for any task $k$ at test time:
$R_{k}(L(S_{u},\mathcal{C}))=\sum_{j=1}^{p}w_{k}^{j}~{}U_{j}(f_{\bm{\theta}_{k}}(S_{u},\mathcal{C})~{}|~{}\bm{\theta}_{k}=g_{\bm{\xi}}(\bm{w}_{k})).$ (3)
@@
3 PaDiRec: The Proposed Approach
In this section, we provide a detailed description of the proposed approach, PaDiRec. PaDiRec utilizes a conditional generative framework designed to directly learn from the optimized parameters of recommendation models tailored to specific tasks. This pre-training process enables the generation of new model parameters based on specified preference weights at test time.
3.1 Algorithm Overview
As shown in Figure 1, we provide an illustrative overview of the proposed PaDiRec, which contains the following three phases. (1) Preparation of adapters: the left part in Figure 1 shows the training process of the recommendation model, from which we can obtain a collection of optimized adapter parameters for feasible specific task by sampling the preference weights. Note that we focus on two utilities: accuracy and diversity. As defined in Eq. (3), each task is represented by a set of preference weights for these two utilities. (2) Parameter diffusion model training: the middle part in Figure 1 illustrates the conditional training procedure of the generative model $g_{\bm{\xi}}$ (i.e., DiT) with the optimized adapter parameters as initial data and the corresponding preference weights as condition, thus generating meaningful adapter parameters from Gaussian noise given preference weights. (3) Test-time parameter generation: the right part in Figure 1 shows how we utilize the trained DiT model during the test phase to adapt to dynamically changing task requirements (i.e., preference weights for diversity and accuracy). First, we quantify these task requirements as preference weights. Next, we employ the trained DiT model to generate adapter parameters in real time, using these preference weights as inputs, which are then combined with the backbone to directly support the recommendation task. In the following subsections, we elaborate on the details of these phases.
Figure 1: An overview of the proposed PaDiRec. Details are shown in Sec. 3.1
3.2 Preparation of Adapters
@@
Model structure. As shown in the left module of Figure 1, sequential recommendation models take user history and candidate items as input. Guided by the objective function (i.e., loss function), the model learns the underlying relationships within the user history, ultimately generating a recommendation list (i.e., Rec. List) from the candidate items. Existing recommender systems based on deep neural networks can be quite large, and making significant invasive modifications typically requires retraining the entire model, which is often prohibitively expensive in industrial applications. To address this, our approach introduces an adapter module, which can be seamlessly integrated into existing sequential recommendation models. Specifically, we incorporate the adapter using a residual connection, attaching it to the last layer of the backbone model. In this setup, the backbone is set to retain the original recommendation capabilities, while the adapter is responsible for adapting to specific tasks.
@@
$\ell_{i}=w_{i}^{1}\ell_{\text{accuracy}}+w_{i}^{2}\ell_{\text{diversity}}.$ (4)
Adapter tuning. Based on above total loss function, we decompose the recommendation model parameters $\bm{\theta}$ into two components: task-specific adapter parameters, denoted as $\bm{\theta}_{\text{a}}$ and task-independent backbone parameters, denoted as $\bm{\theta}_{\text{b}}$. Accordingly, optimizing the model is divided into two phases. The first phase focuses on optimizing the backbone parameters $\bm{\theta}_{\text{b}}$, which uses the standard BCE loss to train the backbone model thus preserving the original recommendation accuracy. The second phase is about the optimization of the task-specific adapter parameters $\bm{\theta}_{\text{a}}$, which aims at improving the system’s adaptability to different tasks. During the second phase, the backbone parameters are frozen to prevent them from being tailored to any specific task, whereas the adapter is trainable. More specifically, in the second phase, we train the task-specific adapter parameters based on two loss functions as in Eq. (4), one for accuracy and one for diversity. For the accuracy loss $\ell_{\text{accuracy}}$, we continue to use BCE as the loss function to guide the model toward accuracy. For the diversity loss $\ell_{\text{diversity}}$, inspired by Yan et al. (2021), we apply a differentiable smoothing of the $\alpha$-DCG metric and adapt it to the recommendation setting. Consider $|\mathcal{C}|$ candidate items and $|\mathcal{M}|$ categories, where each item may cover 0 to $|\mathcal{M}|$ categories. The category labels are denoted as $y_{k,l}$: $y_{k,l}=1$ if item $k$ covers category $m$, and $y_{k,l}=0$ otherwise, where $k\in\{0,\ldots,|\mathcal{C}|-1\}$, $l\in\{0,\ldots,|\mathcal{M}|-1\}$. Based on the $\alpha$-DCG, we design a differentiable diversity loss function:
$\ell_{\text{diversity}}=-\sum_{k=1}^{|\mathcal{C}|}\sum_{l=1}^{|\mathcal{M}|}\frac{y_{k,l}(1-\alpha)C_{k,l}}{\log_{2}(1+\text{Rank}_{k})},$ (5)
@@
$\text{Rank}_{k}=1+\sum_{j\neq k}\text{sigmoid}\left((s_{j}-s_{k})/T\right),\quad C_{k,l}=\sum_{j\neq k}y_{j,l}\cdot\text{sigmoid}\left((s_{j}-s_{k})/T\right),$ (6)
where $s_{k}$ denotes the relevance score of the $k$-th candidate item output by the model. For task $i$, we denote $\bm{\theta}_{i}$ as the model parameters including task-specific adapter parameters $\bm{\theta}_{i}^{\text{a}}$ and fixed backbone parameters $\bm{\theta}_{i}^{\text{b}}$. Overall, based on the total loss in Eq. (4), the task-specific optimization process of $\bm{\theta}_{i}^{\text{a}}$ for task $i$ can be formulated as follows:
$\bm{\theta}_{i}^{\text{a}}=\operatornamewithlimits{arg\,min}_{\bm{\theta}^{\text{a}}_{i}}~{}w_{i}^{1}\ell_{\text{accuracy}}+w_{i}^{2}\ell_{\text{diversity}},$ (7)
@@
3.3 Parameter Diffusion Model Training

### v2

# Paragon: Parameter Generation for Controllable Multi-Task Recommendation

@@ Abstract.
Commercial recommender systems face the challenge that task requirements from platforms or users often change dynamically (e.g., varying preferences for accuracy or diversity). Ideally, the model should be re-trained after resetting a new objective function, adapting to these changes in task requirements. However, in practice, the high computational costs associated with retraining make this process impractical for models already deployed to online environments. This raises a new challenging problem: how to efficiently adapt the learned model to different task requirements by controlling the model parameters after deployment, without the need for retraining. To address this issue, we propose a novel controllable learning approach via parameter generation for controllable multi-task recommendation (Paragon), which allows the customization and adaptation of recommendation model parameters to new task requirements without retraining. Specifically, we first obtain the optimized model parameters through adapter tunning based on the feasible task requirements. Then, we utilize the generative model as a parameter generator, employing classifier-free guidance in conditional training to learn the distribution of optimized model parameters under various task requirements. Finally, the parameter generator is applied to effectively generate model parameters in a test-time adaptation manner given task requirements. Moreover, Paragon seamlessly integrates with various existing recommendation models to enhance their controllability. Extensive experiments on two public datasets and one commercial dataset demonstrate that Paragon can efficiently generate model parameters instead of retraining, reducing computational time by at least 94.6%. The code is released at https://github.com/bubble65/Paragon.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06journalyear: 2025copyright: acmlicensedconference: Proceedings of the Nineteenth ACM Conference on Recommender Systems; September 22–26, 2025; Prague, Czech Republicbooktitle: Proceedings of the Nineteenth ACM Conference on Recommender Systems (RecSys ’25), September 22–26, 2025, Prague, Czech Republicdoi: 10.1145/3705328.3748069isbn: 979-8-4007-1364-4/2025/09ccs: Information systems Recommender systems
(a) Multi-task recommendation model in static environment.
(b) Multi-task recommendation model in dynamic environment.
Figure 1. Illustration of multi-task recommendation in static and dynamic environments. “Aspects” represents the aspects of recommended results such as diversity, fairness, etc. “Preference weights” is a vector of preferences across different aspects. Each new task requirement can be represented by a vector of novel preference weights (maybe different).
1. Introduction
Traditional recommender systems focus on improving accuracy through analyzing user behaviors and contextual data (Kang and McAuley, 2018; Hidasi et al., 2016; Zhang et al., 2024c, b, 2025b, a, 2025a, d). Nowadays, recommendation, especially multi-task recommendation (MTR), places greater emphasis on multiple aspects like diversity and fairness (Xia et al., 2017; Oosterhuis, 2021). They simultaneously optimize multiple aspects based on the task requirements (i.e., preference weights for each aspect). This has made MTR a research hotspot by leveraging unified models to learn interrelated aspects for mutual improvement (He et al., 2022; Yang et al., 2023; Su et al., 2024; Wang et al., 2023a).
Nonetheless, in the online stage of recommendation, there is often a need for sudden, instantaneous changes in task requirements, where the preference weights for each aspect can change unexpectedly. From a commercial perspective, businesses often need real-time adjustments to their recommendation strategies, especially during live events like Black Friday, where user interests can shift drastically in a matter of hours. From a user’s perspective, preferences can change unexpectedly as well. For instance, a user might initially prioritize highly accurate recommendations when searching for a product, but after several similar items, they may start preferring more diverse suggestions to explore new options. These changes are sudden and instantaneous during the online testing stage, making it impossible to predict them in a data-driven manner, highlighting the importance of considering sudden changes in task requirements in MTR to avoid sub-optimal recommendations.
Figure 2. Illustration of controllable multi- task recommendation (CMTR) given task requirement $i$ at testing time.
Traditional MTR algorithms face challenges in addressing this issue (i.e., the sudden changing preference weights on each aspect). Most existing MTR algorithms focus on achieving mutual improvement in static environment. As shown in Figure 1 (a), the task requirement (the preference weights on each aspect such as accuracy and diversity) are predefined and fixed during both training and testing (Zhang and Yang, 2021; Sener and Koltun, 2018). When traditional MTR algorithms are deployed in the dynamic environment, they require reconstructing the objective function, retraining the model, and redeploying it to the online system whenever a new task requirement arises, as depicted in Figure 1 (b). Obviously, when faced with a sudden change in task requirements, it cannot respond immediately due to the highly time- and resource-intensive nature of retraining. Overall, the central issue is that task requirements are sudden changing and manually specified at test time, resulting in unavoidable retraining overhead for new demands. Consequently, developing efficient methods to adjust models to real-time task requirements is crucial, which we term ̵controllable multi-task recommendation (CMTR).
CMTR aims to develop an efficient test-time model control mechanism that eliminates the need for retraining, as illustrated in Figure 2. In fact, retraining essentially involves adjustments to the model parameters. To capture these global parameter changes, we innovatively use a generative model to generate recommendation model parameters instead of relying on the time-consuming retraining process. Additionally, to preserve the foundational capabilities of the recommendation model and reduce the learning complexity and inference time of the generative model, we generate only task-specific adapters. Moreover, treating the dynamically changing task requirements as conditioning signals further confers test‑time controllability on CMTR, thereby eliminating the need to retrain the recommendation model during inference. Specifically, we proposed Paragon, begins by constructing an objective function aligned with task-specific preference weights, and through advanced optimization techniques, we fine-tune recommendation model parameters using adapter tuning. We then train a generative model (e.g., diffusion model) to learn the conditional distribution of these optimized adapter parameters under various task requirements, where the classifier-free guidance training strategy is employed to perform conditional training. Once trained, during online testing, the generative model can rapidly generate task-specific adapter parameters conditioned on the task requirements, which are then integrated with the backbone to produce recommendation lists that meet the specified requirements. Extensive experiments demonstrate that: (1) Paragon can rapidly and instantly generate high-performance model parameters without retraining. (2) The parameters generated by Paragon exhibit strong robustness. (3) Paragon performs well in controlling multiple aspects beyond accuracy and diversity.
We summarize our contributions as follows:
  • We reveal the impact of sudden changes in task requirements during online testing on MTR and define the CMTR task, which emphasizes the model’s adaptability to sudden changing task requirements in online scenarios.
  • We introduce Paragon, which generates recommendation model parameters based on task requirements, eliminating costly retraining.
  • Extensive experiments show that Paragon cuts computational time by at least 94.6% while preserving virtually the same recommendation performance as retraining.
2. Related Work
Multi-task learning (MTL) aims to develop unified models that tackle multiple learning tasks simultaneously while facilitating information sharing (Zhang and Yang, 2021; Ruder, 2017; Shi et al., 2024, 2025; Shen et al., 2024). Recent advancements in MTL include deep networks with various parameter sharing mechanisms (Misra et al., 2016; Long et al., 2017; Yang and Hospedales, 2016) and approaches treating MTL as a multi-objective optimization problem (Lin et al., 2019; Mahapatra and Rajan, 2020; Xie et al., 2021). These latter methods focus on identifying Pareto-efficient solutions across tasks, with significant applications in recommender systems (Jannach, 2022; Li et al., 2020a; Zheng and Wang, 2022). Researchers have explored different strategies, from alternating optimization of joint loss and individual task weights to framing the process as a reinforcement learning problem (Xie et al., 2021). The emphasis has shifted from optimizing specific preference weights to finding weights that achieve Pareto efficiency across objectives (Sener and Koltun, 2018; Lin et al., 2019; Liu et al., 2019). Recently approaches, such as the CMR (Chen et al., 2023a), utilize hypernetworks to learn the trade-off curve for MTL problems. However, our novel approach diverges from these existing methods by employing diffusion models to control model parameters at test time, potentially offering greater flexibility and adaptability in handling multi-task learning problems. Recently, some studies have utilized large language models (LLMs) to influence recommendation systems (Qin et al., 2025a, b) or act as recommenders (Qin et al., 2024b; Xu et al., 2025). However, these methods are unable to provide flexible parameter control and instead depend on semantic understanding, a characteristic that fundamentally differentiates them from our approach. Moreover, there has been a growing body of research focusing on dynamic model adaptation at test time (Shen et al., 2023; Chen et al., 2023a; Shen et al., 2024).
Diffusion models. Diffusion probabilistic models (Ho et al., 2020; Song et al., 2020; Nichol and Dhariwal, 2021)have not only achieved significant success in the field of image generation but have also found wide applications in various other areas in recent years, such as video generation  (Ho et al., 2022b), text generation (Li et al., 2022; Gong et al., 2022), etc. Moreover, diffusion models have shown the ability to generate high-quality neural network parameters, achieving comparable or even superior performance to traditionally trained models (Yuan et al., 2024; Schürholt et al., 2022; Knyazev et al., 2021; Wang et al., 2024). These models have also been applied to enhance the accuracy of recommender systems by addressing challenges such as noisy interactions and temporal shifts in user preferences (Wang et al., 2023c). In our work, we utilize diffusion models to generate parameters for controllable multi-task recommender systems. Notably, we use generative models to generate high-quality parameters for recommendation models. The parameter generation paradigm differs from generative retrieval (Qin et al., 2024a, 2025c), which directly generates item identifiers. Our approach dynamically reconfigures models, maintaining efficiency and enabling control.
3. Problem Formulation and Analyses
@@
To efficiently and effectively adapt to changes in tasks, this paper focuses on controlling the model parameters $\bm{\theta}$ of the recommendation model $f_{\bm{\theta}}$ to accommodate the varying preference weights of new tasks. Specifically, we treat the preference weights as variables and model the relationship between the preference weight space $\mathcal{W}$ and the model parameter space $\Theta$ during training, transforming the time- and resource-intensive retraining problem at test time into an efficient inference problem. Formally, we aim to find a function $g_{\bm{\xi}}:\mathcal{W}\rightarrow\Theta$ (where $\bm{\xi}$ denotes the parameter of $g$) that generates model parameters capable of achieving a high reward given the new preference weights $\bm{w}_{k}$ for any task $k$ at test time:
(3) $R_{k}(L(S_{u},\mathcal{C}))=\sum_{j=1}^{p}w_{k}^{j}~U_{j}(f_{\bm{\theta}_{k}}(S_{u},\mathcal{C})~|~\bm{\theta}_{k}=g_{\bm{\xi}}(\bm{w}_{k})).$
@@
Figure 3. An overview of the proposed Paragon.
[…]
An overview of the proposed Paragon.
4. Paragon: The Proposed Approach
In this section, we provide a detailed description of the proposed approach, Paragon. As shown in Figure 3, we provide an illustrative overview of the proposed Paragon, which contains the following three phases. (1) Preparation of adapters: the left part in Figure 3 shows the training process of the recommendation model, from which we can obtain a collection of optimized adapter parameters for feasible specific task by sampling the preference weights. We focus on two utilities: accuracy and diversity. As defined in Eq. (3), each task is represented by a set of preference weights for these two utilities.(2) Parameter diffusion model training: the middle part in Figure 3 illustrates the conditional training procedure of the generative model $g_{\bm{\xi}}$ (i.e., DiT) with the optimized adapter parameters as initial data and the corresponding preference weights as condition, thus generating meaningful adapter parameters from Gaussian noise given preference weights. (3) Test-time parameter generation: the right part in Figure 3 shows how we utilize the trained DiT model at test time to adapt to dynamically changing task requirements (i.e., preference weights for diversity and accuracy). First, we quantify these task requirements as preference weights. Next, we employ the trained DiT model to generate adapter parameters at test time, using these preference weights as inputs, which are then combined with the backbone to directly support the recommendation task.
4.1. Preparation of Adapters
@@
Model structure. As shown in the left module of Figure 3, sequential recommendation models take user history and candidate items as input. Guided by the objective function (i.e., loss function), the model learns the underlying relationships within the user history, ultimately generating a recommendation list (i.e., Rec. List) from the candidate items. To preserve the foundational capabilities of the recommendation model and reduce the learning complexity and inference time of the generative model. we adopt a backbone recommendation model and task-specific adapters framework, where the backbone is trained to preserve the accurate recommendation by BPR loss while the adapters are trained under various task requirements for generative model to learn. Specifically, we incorporate the adapter using a residual connection, attaching it to the last layer of the backbone model. Based on this structure, the backbone is fixed after training, and on this foundation, the task-specific adapters are independently trained under the guidance of their respective objective functions.
@@
(4) $\ell_{i}=w_{i}^{1}\ell_{\text{accuracy}}+w_{i}^{2}\ell_{\text{diversity}},$
where $\ell_{\text{accuracy}}$ employs the BPR loss (Rendle et al., 2012), and the formulation of $\ell_{\text{diversity}}$ is detailed in the next section.
Adapter tuning. Based on above total loss function, we decompose the recommendation model parameters $\bm{\theta}$ into two components: task-specific adapter parameters, denoted as $\bm{\theta}_{\text{a}}$ and task-independent backbone parameters, denoted as $\bm{\theta}_{\text{b}}$. Accordingly, optimizing the model is divided into two phases. The first phase focuses on optimizing the backbone parameters $\bm{\theta}_{\text{b}}$, which uses the standard BPR loss to train the backbone model thus preserving the original recommendation accuracy. The second phase is about the optimization of the task-specific adapter parameters $\bm{\theta}_{\text{a}}$, which aims at improving the system’s adaptability to different tasks. During the second phase, the backbone parameters are frozen to prevent them from being tailored to any specific task, whereas the adapter is trainable. More specifically, in the second phase, we train the task-specific adapter parameters based on two loss functions as in Eq. (4), one for accuracy and one for diversity. For the accuracy loss $\ell_{\text{accuracy}}$, we continue to use BPR as the loss function to guide the model toward accuracy. For the diversity loss $\ell_{\text{diversity}}$, inspired by Yan et al. (2021), we apply a differentiable smoothing of the $\alpha$-DCG metric and adapt it to the recommendation setting. Consider $|\mathcal{C}|$ candidate items and $|\mathcal{M}|$ categories, where each item may cover 0 to $|\mathcal{M}|$ categories. The category labels are denoted as $y_{k,l}$: $y_{k,l}=1$ if item $k$ covers category $m$, and $y_{k,l}=0$ otherwise, where $k\in\{0,\ldots,|\mathcal{C}|-1\}$, $l\in\{0,\ldots,|\mathcal{M}|-1\}$. Based on the $\alpha$-DCG, we design a differentiable diversity loss function:
(5) $\ell_{\text{diversity}}=-\sum_{k=1}^{|\mathcal{C}|}\sum_{l=1}^{|\mathcal{M}|}\frac{y_{k,l}(1-\alpha)C_{k,l}}{\log_{2}(1+\text{Rank}_{k})},$
@@
(6) $\text{Rank}_{k}$ $=1+\sum_{j\neq k}\text{sigmoid}\left((s_{j}-s_{k})/T\right),$
$\quad C_{k,l}$ $=\sum_{j\neq k}y_{j,l}\cdot\text{sigmoid}\left((s_{j}-s_{k})/T\right),$
where $s_{k}$ denotes the relevance score of the $k$-th candidate item output by the model. For task $i$, we denote $\bm{\theta}_{i}$ as the model parameters including task-specific adapter parameters $\bm{\theta}_{i}^{\text{a}}$ and fixed backbone parameters $\bm{\theta}_{i}^{\text{b}}$. Based on the total loss in Eq. (4), the task-specific optimization process of $\bm{\theta}_{i}^{\text{a}}$ for task $i$ can be formulated as follows:
(7) $\bm{\theta}_{i}^{\text{a}}=\operatornamewithlimits{arg\,min}_{\bm{\theta}^{\text{a}}_{i}}~w_{i}^{1}\ell_{\text{accuracy}}+w_{i}^{2}\ell_{\text{diversity}},$
@@
4.2. Training of Parameter Generation Model